Weighted Average Cost Method with Perpetual Inventory The beginning inventory for Midnight Supplies and data on...

  1. Weighted Average Cost Method with Perpetual Inventory

    The beginning inventory for Midnight Supplies and data on purchases and sales for a three-month period are as follows:

    Date Transaction Number
    of Units
    Per Unit Total
    Jan. 1 Inventory 7,700 $79.00 $608,300
    10 Purchase 23,100 89.00 2,055,900
    28 Sale 11,550 158.00 1,824,900
    30 Sale 3,850 158.00 608,300
    Feb. 5 Sale 1,540 158.00 243,320
    10 Purchase 55,440 91.50 5,072,760
    16 Sale 27,720 168.00 4,656,960
    28 Sale 26,180 168.00 4,398,240
    Mar. 5 Purchase 46,200 93.50 4,319,700
    14 Sale 30,800 168.00 5,174,400
    25 Purchase 7,700 94.00 723,800
    30 Sale 26,950 168.00 4,527,600

    Required:

    1. Record the inventory, purchases, and cost of goods sold data in a perpetual inventory record similar to the one illustrated in Exhibit 5, using the weighted average cost method. Round unit cost to two decimal places, if necessary. Round all total cost amounts to the nearest dollar.

    Midnight Supplies
    Schedule of Cost of Goods Sold
    Weighted Average Cost Method
    For the Three Months Ended March 31
    Purchases Cost of Goods Sold Inventory
    Date Quantity Unit Cost Total Cost Quantity Unit Cost Total Cost Quantity Unit Cost Total Cost
    Jan. 1 $ $
    Jan. 10 $ $
    Jan. 28 $ $
    Jan. 30
    Feb. 5
    Feb. 10
    Feb. 16
    Feb. 28
    Mar. 5
    Mar. 14
    Mar. 25
    Mar. 30
    Mar. 31 Balances $ $

    2. Determine the total sales, the total cost of goods sold, and the gross profit from sales for the period.

    Total sales $
    Total cost of goods sold $
    Gross profit $

    3. Determine the ending inventory cost as of March 31.
    $
Answer #1

1.

Record the inventory, purchases, and cost of goods sold data in a perpetual inventory as follows:

Midnight Supplies
Schedule of Cost of Goods Sold
Weighted Average Cost Method
For the Three Months Ended March 31
Purchases Cost of Goods Sold Inventory
Date Quantity Unit Cost Total Cost Quantity Unit Cost Total Cost Quantity Unit Cost Total Cost
Jan. 1 - - - - - - 7700 $79 $608,300
Jan. 10 23100 $89 $2,055,900 - - - (23100+7700 = 30800) [($608300+2055900)/30800 = $86.5] $2,664,200
Jan. 28 - - - 1150 $86.50 $99,475.00 [30800-1150 = 29650] $86.50 $2,564,725
Jan. 30 - - - 3850 $86.50 $333,025.00 [29650-3850 = 25800] $86.50 $2,231,700
Feb. 5 - - - 1540 $86.50 $133,210.00 [25800-1540 = 24260] $86.50 $2,098,490
Feb. 10 55440 $91.50 $5,072,760 - - - [24260+55440 = 79700] [($2098490+$5072760)/79700 = $89.98] $7,171,406
Feb. 16 - - - 27720 $89.98 $2,494,246 [79700-27720 = 51980] $89.98 $4,677,160
Feb. 28 - - - 26180 $89.98 $2,355,676 [51980-26180 = 25800] $89.98 $2,321,484
Mar. 5 46200 $93.50 $4,319,700 [25800+46200 = 72000] [($2321484+$4319700)/72000 = $92.23] $6,640,560
Mar. 14 30800 $92.23 $2,840,684 [72000-30800 = 41200] $92.23 $3,799,876
Mar. 25 7700 $94 $723,800 [41200+7700 = 48900] [($3799876+$723800)/48900 = $92.50] $4,523,250
Mar. 30 26950 $92.50 $2,492,875 [48900-26950 = 21950] $92.50 $2,030,375
Mar. 31 Balances 118190 $10,749,191.00 21950 $92.50 $2,030,375

_____________________________________________________________________________________

2.

Determine the total sales, the total cost of goods sold, and the gross profit from sales for the period.

Quantity Sale price Total sales
1150 $158 $181,700
3850 $158 $608,300
1540 $158 $243,320
27720 $168 $4,656,960
26180 $168 $4,398,240
30800 $168 $5,174,400
26950 $168 $4,527,600
Total $19,790,520
Total Sales $19,790,520
Less: COGS ($10,749,191)
Gross profit $9,041,329

_____________________________________________________

3.

Ending inventory cost as of March 31 is $2030375 (see part 1)